OpenStudy (moongazer):

If a ladder 13m long leaning against a wall slides 1m down the wall, how long will be the new base if the previous base was 5m long?

OpenStudy (anonymous):

H_old=sqrt(13^2-5^2)=sqrt(169-25)=sqrt(144)=12 H_new=H_old-1=11 B_new=sqrt(13^2-11^2)=sqrt(169-121)=sqrt(48)=4sqrt(3)~6.85